OpenStudy (anonymous):

Express the radical in simplest form: (square root) 1/11

OpenStudy (anonymous):

@amoodarya

OpenStudy (amoodarya):

let's do it together

OpenStudy (amoodarya):

multiply by 11/11

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

mult. 11x11?

OpenStudy (anonymous):

hello?

OpenStudy (amoodarya):

\[\sqrt{\frac{ 1 }{ 11 }}=\\ \sqrt{\frac{ 1 }{ 11 }\frac{ 11 }{ 11 }}=\\ \sqrt{\frac{ 11 }{ 11^2 }}=\frac{\sqrt{11}}{11}\]